# Homeomorphisms of continua through projective Fraïssé limits

### v1

@@ 1. $\boxdot_{6}$
    $f_{k}(a)=l_{\pi_{4k+2,4k+3}(a)+2},$
or for almost all $a$’s $f_{k}(a)=l_{\pi_{4k+2,4k+3}(a)-2}$ holds. We will assume $\boxdot_{6}$, as the other case is essentially the same. We
@@ 1. $\boxdot_{7}$
holds by $\boxdot_{5}$ (since the $W_{j}$’s are pairwise disjoint). Pick $\bar{z}=(z_{i})_{i}\in W_{2}\cap\pi_{4k+2}^{-1}(l_{m})$. Now $\pi_{4k+3,4k+2}(z_{4k+3})=z_{4k+2}=l_{m}$, so necessarily
@@
Recalling (5.8) and the fact that $\sigma({\rm pr}(\bar{z}))$ is the limit of ${\rm pr}(\sigma_{i}(\bar{z}))$ (in $P$) by compactness of $\mathbb{P}$} there exists a sequence $\bar{z}^{*}=(z^{*}_{i})_{i}\in{\mathbb{P}}$ that is a an accumulation point of $(\sigma_{i}(\bar{z}))_{i}$, so necessarily represents $\sigma({\rm pr}(\bar{z}))\in P$, i.e. ${\rm pr}(\bar{z})=\sigma({\rm pr}(\bar{z}))$ and has the property
$z^{*}_{4k+3}\ R\ (\pi_{4k+3}\circ\sigma_{k+1}(\bar{z}))=h_{k+1}(z_{4k+4})$
(where the equality follows from the way we picked $\sigma_{k}$, $(4)$). This means that the following claim implies $(z^{*}_{i})_{i}\in W_{3}$, so ${\rm pr}((z^{*}_{i})_{i})\in U_{3}$, and we are done.
@@
If $b\in J_{4k+3}$ is such that $bRh_{k+1}(z_{4k+4})$, then $\pi_{4k+3}^{-1}(b)\subseteq W_{3}$.
@@
First we let $w=h_{k+1}(z_{4k+4})\in L_{4k+3}$, note that (by ($\blacksquare_{7}$$\blacksquare_{7}(b)$))
$\pi_{4k+2,4k+3}(w)=\pi_{4k+2,4k+3}(h_{k+1}(z_{4k+4}))=f_{k}(z_{4k+3})=l_{m+2},$
in particular, $\neg(\pi_{4k+2,4k+3}(w)Rl_{m}$, so by $\blacksquare_{6}$ and (5.18)
(5.20) $wR^{3}b\ \Rightarrow\ \pi_{4k+2,4k+3}(b)\in\{l_{m+1},l_{m+2},l_{m+3}\}\subseteq L_{4k+2}.$
This means $bR^{3}w$ implies $\pi_{4k+3}^{-1}(b)\cap W_{2}=\emptyset$ by the definition of $m$ ($\boxdot_{7}$) Hence, by the definition of the $W_{j}$’s (i.e. $W_{i}={\rm pr}^{-1}(U_{i})$)
@@
To summarize, we got
$\bigcup_{b\in J_{4k+3},\ bR^{3}w}{\rm pr}(\pi_{4k+3}^{-1}(b))\cap(U_{1}\cup U_{2})=\emptyset.$
Since $\bigcup_{j=1}^{3}U_{j}$ is dense in $P$, $U_{3}$ must be dense in the open set
$P\setminus\left(\bigcup_{b\in J_{4k+3},\ \neg(bR^{3}w)}{\rm pr}(\pi_{4k+3}^{-1}(b))\right),$
so using that ${\rm pr}(\pi_{4k+3}^{-1}(d))$, ${\rm pr}(\pi_{4k+3}^{-1}(c))$ are closed sets that are from positive distance apart iff $\neg(cRd)$,
$\bigcup_{b\in J_{4k+3},\ bR^{2}w}{\rm pr}(\pi_{4k+3}^{-1}(b))\subseteq P\setminus\left(\bigcup_{b\in J_{4k+3},\ \neg(bR^{3}w)}{\rm pr}(\pi_{4k+3}^{-1}(b))\right)\subseteq$
$\subseteq{\rm int}({\rm cl}(U_{3}))={\rm int}(U_{3})=U_{3},$
as $U_{3}$ is a regular open set, which finishes the proof of the lemma.
The inductive steps (depending on the remainder of $i$ modulo $4$) will rely on Lemmas 5.11-5.14 below, with Lemma 5.11 handling the case of stepping from $4k$ to $4k+1$ (constructing $J_{4k+1}$, $\pi_{4k,4k+1}$, $h^{\bullet}_{k}$), and the last one handling the case of going from $4k+3$ to $4k+4$ (constructing $J_{4k+4}$, $\pi_{4k+3,4k+4}$, $h_{k+1}$). Before stating these lemmas we first prove some auxiliary lemmas about piecewise amalgamation with some values prescribed.
@@
    for $c\in C^{\prime}_{1}\cup C^{\prime}_{2}$, if $vRK^{\prime}$ (in particular, $c$ is an endpoint of $C_{j}^{\prime}$ for $j=1$ or $j=2$), then
@@ 6. $\boxplus_{6}$
Now let $J^{+}$ be a linear graph which is the disjoint union of the intervals $C_{1}^{+}$, $K^{+}$, $C_{2}^{+}$ (where $K^{+}$ lies in the middle). Pick a finite linear graph $J^{\prime\prime}$ which is the disjoint union of the intervals $C_{1}^{\prime\prime}$, $K^{\prime\prime}$, $C_{2}^{\prime\prime}$ and $\pi^{*}\in{\rm Epi}(J^{\prime\prime},J^{+})$ such that $\pi^{*}[K^{\prime\prime}]=K^{+}$, $\pi^{*}[C_{i}^{\prime\prime}]=J^{+}$, moreover,
@@
    $\pi^{*}(a_{0}),\pi^{*}(a_{1}),\ldots,\pi^{*}(a_{|C^{\prime\prime}_{1}|-1})$
    visits all nodes of $C^{+}_{1}$ before reaching $K^{+}$, i.e.
    ${\rm tp}^{a_{0},C_{1}^{\prime\prime}}(\pi^{*})\supseteq\{C^{+}_{1},C^{+}_{1}\cup K^{+}\},$
@@
    ${\rm tp}^{b_{0},C_{2}^{\prime\prime}}(\pi^{*})\supseteq\{C^{+}_{2},C^{+}_{2}\cup K^{+}\},$
  • for every $d\in J^{+}$, each connected component of $(\pi^{*})^{-1}(d)$ is an interval including at least $3$ nodes.
One easily checks that $\pi^{\prime}=\pi^{*}\circ\pi^{+}$, $h^{\bullet}=\pi^{*}\circ h^{+}$ will work.
@@ 7. $\boxplus_{7}$
Notice that the main difference compared to Lemma 5.11 is that we “double” $K^{\prime\prime}$ (and we have $3$ complementer intervals instead of $2$). Let $J^{*}$, $K^{*}$, $C_{1}^{*}$, $C_{2}^{*}$, $\pi^{*}$, $g^{*}$ be $J^{\prime\prime}$, $K^{\prime\prime}$, $C_{1}^{\prime\prime}$, $C_{2}^{\prime\prime}$, $\pi^{\prime\prime}$ and $h^{\bullet}$ given by Lemma 5.11. Our $J^{\prime\prime}$ will be an appropriate extension of $J^{+}$.
So let $J^{\prime\prime}$ be a finite linear graph with an epimorphism $\pi^{**}:J^{\prime\prime}\to J^{*}$ that satisfies
@@
  • $\pi^{**}[C^{\prime\prime}_{1}]=\pi^{**}[C^{\prime\prime}_{3}]=C^{*}_{1}$,
  • $\pi^{**}[C^{\prime\prime}_{2}]=C_{2}^{*}$,
  • $\pi^{**}[K^{\prime\prime}]=\pi^{**}[L^{\prime\prime}]=K^{*}$,
  • for every $d\in J^{*}$, each connected component of $(\pi^{**})^{-1}(d)$ is an interval including at least $3$ nodes.
It is straightforward to check that $\pi^{\prime}=\pi^{*}\circ\pi^{**}$, $g=g^{*}\circ\pi^{**}$ work. ∎
@@
    $\pi[L^{\prime}]=K$ ($g[K^{\prime}]$), $\pi$ maps endpoints to endpoints,
@@ 8. $\boxplus_{8}$
    for each $a\in L^{\prime\prime}$ we have $|f(a)-f^{\prime}(a)|\leq 2$, and if $a\in L^{\prime\prime}$ is such that $f^{\prime}(a)=\pi^{\prime}(a)$ is neither an endpoint of $L^{\prime}$ nor is related to one then $|f(a)-f^{\prime}(a)|=2$,
@@
separately, in such a way that they map endpoints to endpoints. If these restrictions satisfy $\boxplus_{1}$, then we will have $\boxplus_{1}$-$\boxplus_{5}$.
We can apply the moreover part of Lemma B.1 separately to the pairs $g{\restriction}C^{\prime}_{i}$ and $\pi{\restriction}C^{\prime}_{i}$ (so $M=M^{\prime}=C^{\prime}_{i}$, $f=g{\restriction}C^{\prime}_{i}$, $f^{\prime}=\pi{\restriction}C^{\prime}_{i}$) where $i=1,3$, to obtain $C_{i}^{\prime\prime}$, $f^{\prime}{\restriction}C_{i}^{\prime\prime}$, $\pi^{\prime}{\restriction}C_{i}^{\prime\prime}$.
Next we apply the main part of Lemma B.1 to $C_{2}$. Finally, since $\pi{\restriction}K^{\prime}=g{\restriction}K^{\prime}$, and $\pi{\restriction}L^{\prime}=g{\restriction}L^{\prime}$ we can take $K^{\prime\prime}=K^{\prime}$, $L^{\prime\prime}=L^{\prime}$ (and $f^{\prime}{\restriction}K^{\prime\prime}=\pi^{\prime}{\restriction}K^{\prime\prime}={\rm id}_{K^{\prime\prime}}$).
@@
Finally, to get $f$ we first note that $(f^{\prime})^{-1}(L^{\prime})=L^{\prime\prime}$ implies that the endpoints of $L^{\prime\prime}$ are sent to those of $L^{\prime}$, i.e. if $l^{\prime\prime}_{-},l^{\prime\prime}_{+}\in L^{\prime\prime}$ are such that $l^{\prime\prime}_{-}RC_{1}^{\prime\prime}$, $l^{\prime\prime}_{+}RC_{2}^{\prime\prime}$, and the endpoints $l^{\prime}_{-},l^{\prime}_{+}\in L^{\prime}$ are such that $l^{\prime}_{-}RC_{1}^{\prime}$, $l^{\prime}_{+}RC_{2}^{\prime}$, then $f^{\prime}(l^{\prime\prime}_{-})=l^{\prime}_{-}$, $f^{\prime}(l^{\prime\prime}_{+})=l^{\prime}_{+}$ (for this we used also $f^{\prime}[C_{i}^{\prime\prime}]=C_{i}^{\prime}$). We also note that by $\boxplus_{6}$ there is a connected subgraph of $L^{\prime\prime}$ containing $l^{\prime\prime}_{-}$, consisting of at least $3$ elements all of which are mapped to $l^{\prime}_{-}$ by $f^{\prime}$.
@@ 4. $\boxplus_{4}$
    $\pi^{\prime}{\restriction}K^{\prime\prime}=h{\restriction}K^{\prime\prime}$,
@@ 7. $\boxplus_{7}$
  • $J^{*}$ and $\pi^{*},h^{*}\in{\rm Epi}(J^{*},J^{\prime})$ with $f\circ\pi^{\prime}=\pi\circ h^{*}$,
  • $J^{**}$ and $\pi^{**}\in{\rm Epi}(J^{**},J^{*})$,
  • and finally $J^{\prime\prime}$, $\pi^{\prime\prime}\in{\rm Epi}(J^{\prime\prime},J^{*})$,
and we will let $\pi^{\prime}=\pi^{*}\circ\pi^{**}\circ\pi^{\prime\prime}$, $h=h^{*}\circ\pi^{**}\circ\pi^{\prime\prime}$.
The extension $J^{*}$ will be the disjoint union of the finite linear graphs $C_{1}^{*}$, $L^{*}$, $C_{2}^{*}$, $K^{*}$, $C_{3}^{*}$, which are connected to each other in this order. $C_{1}^{*}$, $\pi^{*}{\restriction}C_{1}^{*}$, $h^{*}{\restriction}C_{1}^{*}$ are gotten by applying Lemma B.1 to $g{\restriction}C_{1}$ and $\pi{\restriction}C_{1}$, and similarly, we obtain $D\in\{L^{*},C_{2}^{*},C_{3}^{*}\}$, $\pi^{*}{\restriction}D$, $h^{*}{\restriction}D$ by the same way, while we can let $K^{*}$ be isomorphic to $K^{\prime}$ (since $f{\restriction}K^{\prime}=\pi{\restriction}K^{\prime}$), so

### v2

@@ 1. $\boxdot_{6}$
    $f_{k}(a)=\pi_{4k+2,4k+3}(a)\dotplus 2,$
or the other way around, for all but 44 $a$’s $f_{k}(a)=\pi_{4k+2,4k+3}(a)\dotplus(-2)$ holds. We will assume $\boxdot_{6}$, as the other case is essentially the same. We
@@ 1. $\boxdot_{7}$
holds by $\boxdot_{5}$ (since the $W_{j}$’s are pairwise disjoint). Pick $\bar{z}=(z_{i})_{i}\in W_{2}\cap\pi_{4k+2}^{-1}(l_{m})$ (so ${\rm pr}(\bar{z})\in U_{2}$). In the rest of the proof we will show that $\sigma({\rm pr}(\bar{z}))\in U_{3}$.
Now $\pi_{4k+3,4k+2}(z_{4k+3})=z_{4k+2}=l_{m}$, so necessarily
@@
It is easily seen that $\pi_{4k+2,4k+3}\circ h_{k+1}=f_{k}\circ\pi_{4k+3,4k+4}(z_{4k+4})=l_{m+2}$, so $\sigma_{k+1}(\bar{z})\notin W_{2}$. The purpose of the following claim is to show that not only do we have $\sigma_{k+1}(\bar{z})\notin(W_{1}\cup W_{2})$, but points close enough to $\sigma_{k+1}(\bar{z})$ lie outside $W_{1}\cup W_{2}$.
@@
If $b\in J_{4k+3}$ is such that $bR^{3}h_{k+1}(z_{4k+4})$, then $\pi_{4k+3}^{-1}(b)\cap W_{2}=\emptyset$.
@@
By $\boxdot_{7}$ it suffices to show that for any such $b$
$\pi_{4k+2,4k+3}(b)=l_{j}\ \textrm{ for some }j>m.$
We note that (by ($\blacksquare_{7}$$\blacksquare_{7}(b)$))
$\pi_{4k+2,4k+3}(h_{k+1}(z_{4k+4}))=f_{k}(z_{4k+3})=l_{m+2},$
so by $\blacksquare_{6}$ and (5.29)
(5.31) $h_{k+1}(z_{4k+4})R^{3}b\ \Rightarrow\ \pi_{4k+2,4k+3}(b)\in\{l_{m+1},l_{m+2},l_{m+3}\}\subseteq L_{4k+2}.$
We are going to argue that
(5.32) $(bR^{2}h_{k+1}(z_{4k+4}))\ \Rightarrow\ \pi_{4k+3}^{-1}(b)\subseteq W_{3}$
(in fact we will only use the statement for $b$’s satisfying $bRh_{k+1}(z_{4k+4})$). By the definition of the $W_{j}$’s (that is, $W_{i}={\rm pr}^{-1}(U_{i})$)
@@
Let $B=\{b\in J_{4k+3}:\ bR^{3}h_{k+1}(z_{4k+4})\}$, $A=J_{4k+3}\setminus B$, and let
$D_{X}=\bigcup_{x\in X}{\rm pr}(\pi_{4k+3}^{-1}(x))\ \textrm{ for }X\subseteq J_{4k+3},$
so (5.33), (5.34) can be summarized as
$D_{B}\cap(U_{1}\cup U_{2})=\emptyset.$
Since $\bigcup_{j=1}^{3}U_{j}$ is dense in $P$, and $P\setminus D_{A}\subseteq D_{B}$ is open ((i) of Proposition A.4), $U_{3}$ is dense in $P\setminus D_{A}$. But $U_{3}$ is regular open, so
$P\setminus D_{A}\subseteq{\rm int}({\rm cl}(U_{3}))=U_{3}.$
Now using that $D_{X}\cap D_{Y}=\emptyset$ if $\neg(XRY)$, we obtain
$\textrm{if }bR^{2}h_{k+1}(z_{4k+4}),\textrm{ then }{\rm pr}(\pi_{4k+3}^{-1}(b))\subseteq P\setminus D_{A}\subseteq U_{3},$
which yields (5.32) as desired.
It remains to check that the above statements imply $\sigma({\rm pr}(\bar{z}))\in U_{3}$. Recalling (5.21) and the fact that $\sigma({\rm pr}(\bar{z}))$ is the limit of ${\rm pr}(\sigma_{i}(\bar{z}))$ (in $P$), by compactness of $\mathbb{P}$} there exists a sequence $\bar{z}^{*}=(z^{*}_{i})_{i}\in{\mathbb{P}}$ that is a an accumulation point of $\{(\sigma_{i}(\bar{z})):\ i\in\omega\}$ (in $\mathbb{P}$}), so necessarily represents $\sigma({\rm pr}(\bar{z}))\in P$, that is, ${\rm pr}(\bar{z}^{*})=\sigma({\rm pr}(\bar{z}))$ and has the property
$z^{*}_{4k+3}\ R\ \pi_{4k+3}(\sigma_{k+1}(\bar{z}))=h_{k+1}(z_{4k+4})$
(where the equality follows from the way we picked $\sigma_{k}$, $(4)$). This means that (5.32) implies $\bar{z}^{*}=(z^{*}_{i})_{i}\in W_{3}$, so $\sigma({\rm pr}(\bar{z}))={\rm pr}(\bar{z}^{*})\in U_{3}$, and we are done.
The inductive steps (depending on the remainder of $i$ modulo $4$) will rely on Lemmas 5.10-5.13 below, with Lemma 5.10 handling the case of stepping from $4k$ to $4k+1$ (constructing $J_{4k+1}$, $\pi_{4k,4k+1}$, $h^{\bullet}_{k}$ from the inputs $h_{k},\pi_{4k-1,4k}\in{\rm Epi}(J_{4k},J_{4k-1})$), and the last one handling the case of going from $4k+3$ to $4k+4$ (constructing $J_{4k+4}$, $\pi_{4k+3,4k+4}$, $h_{k+1}$). Each lemma will rely on Lemma B.2.
@@
    for $c\in C^{\prime}_{1}\cup C^{\prime}_{2}$, if $cRK^{\prime}$ (in particular, $c$ must be an endpoint of $C_{j}^{\prime}$ for $j=1$ or $j=2$), then
@@ 6. $\boxplus_{6}$
Now let $J^{+}$ be a linear graph which is the disjoint union of the intervals $C_{1}^{+}$, $K^{+}$, $C_{2}^{+}$ (where $K^{+}$ lies in the middle). Pick a finite linear graph $J^{\prime\prime}$ which is the disjoint union of the intervals $C_{1}^{\prime\prime}$, $K^{\prime\prime}$, $C_{2}^{\prime\prime}$ and $\pi^{++}\in{\rm Epi}(J^{\prime\prime},J^{+})$ such that $\pi^{++}[K^{\prime\prime}]=K^{+}$, $\pi^{++}[C_{i}^{\prime\prime}]=J^{+}$, moreover,
@@
    $\pi^{++}(a_{0}),\pi^{++}(a_{1}),\ldots,\pi^{++}(a_{|C^{\prime\prime}_{1}|-1})$
    visits all nodes of $C^{+}_{1}$ before reaching $K^{+}$, that is,
    ${\rm tp}^{a_{0},C_{1}^{\prime\prime}}(\pi^{++})\supseteq\{C^{+}_{1},C^{+}_{1}\cup K^{+}\},$
@@
    ${\rm tp}^{b_{0},C_{2}^{\prime\prime}}(\pi^{++})\supseteq\{C^{+}_{2},C^{+}_{2}\cup K^{+}\},$
  • for every $d\in J^{+}$, each connected component of $(\pi^{++})^{-1}(d)$ is an interval including at least $3$ nodes.
One easily checks that $\pi^{\prime}=\pi^{+}\circ\pi^{++}$, $h^{\bullet}=h^{+}\circ\pi^{++}$ work.
@@ 7. $\boxplus_{7}$
Notice that the main difference compared to Lemma 5.10 is that we “double” $K^{\prime\prime}$ (and we have $3$ complementer intervals instead of $2$). Let $J^{+}$, $K^{+}$, $C_{1}^{+}$, $C_{2}^{+}$, $\pi^{+}$, $g^{+}$ be $J^{\prime\prime}$, $K^{\prime\prime}$, $C_{1}^{\prime\prime}$, $C_{2}^{\prime\prime}$, $\pi^{\prime\prime}$ and $h^{\bullet}$ given by Lemma 5.10. Our $J^{\prime\prime}$ will be an appropriate extension of $J^{+}$.
So let $J^{\prime\prime}$ be a finite linear graph with an epimorphism $\pi^{++}:J^{\prime\prime}\to J^{+}$ that satisfies
@@
  • $\pi^{++}[C^{\prime\prime}_{1}]=\pi^{++}[C^{\prime\prime}_{3}]=C^{*}_{1}$,
  • $\pi^{++}[C^{\prime\prime}_{2}]=C_{2}^{+}$,
  • $\pi^{++}[K^{\prime\prime}]=\pi^{++}[L^{\prime\prime}]=K^{+}$,
  • for every $d\in J^{+}$, each connected component of $(\pi^{++})^{-1}(d)$ is an interval including at least $3$ nodes.
It is straightforward to check that $\pi^{\prime}=\pi^{+}\circ\pi^{++}$, $g=g^{+}\circ\pi^{++}$ work. ∎
@@
    $\pi[L^{\prime}]=K$ ($g[L^{\prime}]$), $\pi$ maps endpoints to endpoints,
@@ 8. $\boxplus_{8}$
    for each $a\in L^{\prime\prime}$ we have ${\rm dist}(f(a),f^{\prime}(a))\leq 2$, and if $a\in L^{\prime\prime}$ is such that $f^{\prime}(a)=\pi^{\prime}(a)$ is neither an endpoint of $L^{\prime}$ nor is related to one then ${\rm dist}(f(a),f^{\prime}(a))=2$,
@@
separately, in such a way that they map endpoints to endpoints. If these restrictions satisfy $\boxplus_{1}$, then (gluing the graphs in the obvious way) we will automatically have $\boxplus_{1}$-$\boxplus_{5}$.
We can apply the moreover part of Lemma B.2 separately to the pairs $g{\restriction}C^{\prime}_{i}$ and $\pi{\restriction}C^{\prime}_{i}$ where $i=1,3$, to obtain $C_{i}^{\prime\prime}$, $f^{\prime}{\restriction}C_{i}^{\prime\prime}$, $\pi^{\prime}{\restriction}C_{i}^{\prime\prime}$ satisfying $\pi\circ f^{\prime}{\restriction}C_{i}^{\prime\prime}=g\circ\pi^{\prime}{\restriction}C_{i}^{\prime\prime}$.
Next we apply the main part of Lemma B.2 to $C_{2}$.
Finally, since $\pi{\restriction}K^{\prime}=g{\restriction}K^{\prime}$, and $\pi{\restriction}L^{\prime}=g{\restriction}L^{\prime}$ we can take $K^{\prime\prime}=K^{\prime}$, $L^{\prime\prime}=L^{\prime}$ (and $f^{\prime}{\restriction}K^{\prime\prime}=\pi^{\prime}{\restriction}K^{\prime\prime}={\rm id}_{K^{\prime\prime}}$).
@@
Finally, to get $f$ we first note that $(f^{\prime})^{-1}(L^{\prime})=L^{\prime\prime}$ implies that the endpoints of $L^{\prime\prime}$ are sent to those of $L^{\prime}$, that is, if $l^{\prime\prime}_{-},l^{\prime\prime}_{+}\in L^{\prime\prime}$ are such that $l^{\prime\prime}_{-}RC_{1}^{\prime\prime}$, $l^{\prime\prime}_{+}RC_{2}^{\prime\prime}$, and the endpoints $l^{\prime}_{-},l^{\prime}_{+}\in L^{\prime}$ are such that $l^{\prime}_{-}RC_{1}^{\prime}$, $l^{\prime}_{+}RC_{2}^{\prime}$, then $f^{\prime}(l^{\prime\prime}_{-})=l^{\prime}_{-}$, $f^{\prime}(l^{\prime\prime}_{+})=l^{\prime}_{+}$ (for this we used also $f^{\prime}[C_{i}^{\prime\prime}]=C_{i}^{\prime}$). We also note that by $\boxplus_{6}$ there is a connected subgraph of $L^{\prime\prime}$ containing $l^{\prime\prime}_{-}$, consisting of at least $3$ elements all of which are mapped to $l^{\prime}_{-}$ by $f^{\prime}$.
@@ 4. $\boxplus_{4}$
    $\pi^{\prime}{\restriction}K^{\prime\prime}=h{\restriction}K^{\prime\prime}$, mapping endpoints to endpoints,
@@ 7. $\boxplus_{7}$
  • $J^{+}$ and $\pi^{+},h^{+}\in{\rm Epi}(J^{+},J^{\prime})$ with $f\circ\pi^{\prime}=\pi\circ h^{+}$,
  • $J^{++}$ and $\pi^{++}\in{\rm Epi}(J^{++},J^{+})$,
  • and finally $J^{\prime\prime}$, $\pi^{\prime\prime}\in{\rm Epi}(J^{\prime\prime},J^{+})$,
and we will let $\pi^{\prime}=\pi^{+}\circ\pi^{++}\circ\pi^{\prime\prime}$, $h=h^{+}\circ\pi^{++}\circ\pi^{\prime\prime}$.
The extension $J^{+}$ will be the disjoint union of the finite linear graphs $C_{1}^{+}$, $L^{+}$, $C_{2}^{+}$, $K^{+}$, $C_{3}^{+}$, which are connected to each other in this order. $C_{1}^{+}$, $\pi^{+}{\restriction}C_{1}^{+}$, $h^{+}{\restriction}C_{1}^{+}$ are gotten by applying Lemma B.2 to $g{\restriction}C_{1}$ and $\pi{\restriction}C_{1}$, and similarly, we obtain $D\in\{L^{+},C_{2}^{+},C_{3}^{+}\}$, $\pi^{+}{\restriction}D$, $h^{+}{\restriction}D$ by the same way, while we can let $K^{+}$ be isomorphic to $K^{\prime}$ (since $f{\restriction}K^{\prime}=\pi{\restriction}K^{\prime}$), so